# Reconciling Different Theories of Learning with an Agent-based Model of Procedural Learning

### v1

@@ 4.2 Expertise Reversal Effect
Interestingly, Figure 1b also provides support for the expertise reversal effect. We find that when learners have low prior knowledge, the ICAP hypothesis does not hold. In particular, until step 25, we find that passive learning is consistently at least as good as (and typically better than) active and constructive learning. This is consistent with the worked example effect. However, after step 25, active and constructive learning gradually overtake passive learning, consistent with the expertise reversal effect. The reason for this is that early on, the increased cognitive load is too high to make active or constructive learning worthwhile, but as the learners learn more, the support of neighboring nodes (and the reinforcement of edges in the case of constructive learning) makes these kinds of learning activities more beneficial, despite the increased cognitive load. This shows why it could be important to take both cognitive load considerations and ICAP considerations into account to make accurate predictions about learning.
4.3 Replication Experiment
To evaluate the performance of the model and validate it, we need to replicate previous empirical studies that have investigated learning procedural knowledge from the lens of the ICAP framework. McEldoon studied the effect of self-explanations and worked examples in learning one-way and two-way ANOVA [10]. Her study consisted of two groups of less and more focused processing, lessFP and moreFP respectively. In the first group, students had to “only determine missing final values within problem-solving steps” and “receive instructional explanations and then copy or paraphrase them” while the more focused processing group had to “determine missing intermediate values within problem-solving steps” and “generate self-explanations before receiving instructional explanations” [10].
Using our model to try to replicate this work is of particular interest as (1) the intervention was designed from the lens of ICAP (lessFP corresponds to active and moreFP corresponds to constructive learning [10, 2]) but the author looks at procedural as well as conceptual learning, and (2) the study tries to assess the amount of cognitive load students experience using students’ self-reports. Furthermore, this dissertation is cited by Chi and Wylie in comparing CLT with ICAP [2].
Based on this work, students in the constructive group reported significantly higher cognitive load compared to active learners [10]. Furthermore, the study was conducted across two weeks with two different topics (i.e., one-way and two-way ANOVA) and students’ learning gains were reported separately for each week. McEldoon finds a significant difference between the two conditions in terms of procedural knowledge at the end of week 1. However, interestingly, the differences between the moreFP and lessFP groups are no longer significant at the end of week 2. According to the author, the discrepancy between week 1 and week 2 is potentially due to two main reasons. First, as per the expertise reversal effect, worked examples work better for novice learners rather than for learners who already have gained some domain knowledge [6, 7, 5] and as the topics of these two weeks share a lot of similar concepts and procedural calculations, the students in these groups are already familiar with the domain after week 1. This is confirmed by the pretest results before week 2. Second, based on self-reports, students in the moreFP group put less effort in week 2 compared to the prior week.
(a) Replication of Week 1 results
(b) Differences between ideal and observed groups in week 2
Figure 2: Replication of McEldoon’s study [10]
We conduct a simulation of their experiment by taking into consideration these two phenomena (observed condition) and comparing them to a simulation of an idealized version of their experiment (idealized condition). In this simulation, the average number of edges for each node is set to 4 as the KCs in the content in both weeks are quite similar and interrelated. In the idealized condition, we simply model moreFP students as active learners in both weeks and lessFP students as constructive learners in both weeks. In the observed condition, we have the moreFP group change their mode of engagement from constructive to active after week 1; this models the lower degree of effort self-reported by students. To model the fact that the topics in week 1 and week 2 overlap in content (and hence learners have a high degree of prior knowledge), we also have both lessFP and moreFP learners practice 5 nodes (steps 20-24) that they had already practiced in week 1.
We find that our simulation shows a difference in the conditions at the end of week 1 (Figure 2a), comparable to the actual experimental results. Moreover, notice that this is despite the fact that students in the moreFP condition self-reported exerting higher effort on the task (i.e., higher cognitive load). This can be seen in the fact that for the first ten or so steps at the beginning of week 1, there is little-to-no difference between the conditions (Figure 2b). However, Figure 2b also shows that the difference in knowledge between moreFP and lessFP learners at the end of week 2 is virtually the same as at the beginning of week 2. This seems to be largely because in the observed condition, the moreFP learners switch to active learning (identical to the lessFP learners). The simulation, through a simplified version of the empirical study, provides insights into the effectiveness of instructional activities in specific scenarios.
5 Discussion and Future Work
@@
Our model represents a more nuanced version of ICAP that is more aligned with empirical results from learning procedural knowledge. Our results show that by modeling cognitive load, the ICAP framework may be able to robustly explain prior empirical literature beyond results on purely conceptual learning. Here we discuss three directions we could pursue in future work to truly capitalize on the benefits of Procedural ABICAP.
5.1 Replication Studies
@@
While we have presented results replicating some major findings in the literature (e.g., the ICAP hypothesis and the expertise reversal effect) as well as a particular study [10], we note that these replication studies are still preliminary. In future work, we could do more precise replications of prior findings. For example, here we showed how conceptually our model could replicate expertise reversal effects. However, we do not know if it would actually predict expertise reversal effects noticed in the literature. By trying to replicate specific studies that noticed such an effect, we may get better insights into the validity of our model (and how it may need to be adjusted). Moreover, Procedural ABICAP could be used to replicate a variety of other results in the literature as well. For example, we could try to replicate results on studies that interleave worked examples and problem-solving tasks [15].
5.2 Making Novel Predictions
@@
Although we have primarily focused on replicating prior results, the true power of agent-based modeling lies in making predictions about novel settings where there is little empirical literature. This could be used to generate hypotheses which could subsequently be verified with experimental studies [11].
@@
Researchers who design interventions based on the ICAP framework often assume that learning gains from higher modes of engagement are higher [3] but as we have shown with our model, this is not always the case. Using a computational model could help elucidate when we expect certain instructional interventions to be more effective than others. Of course, for the simulations to be accurate, it would help to have a concrete sense of the knowledge graph, the cognitive load for different modes of engagement, etc. If researchers have a lot of uncertainty about certain parameters (e.g., the relative cognitive load for different activities), they can use robustness tests to assess the sensitivity of the simulation results to the choice of those parameters. Moreover, models like Procedural ABICAP could also support researchers who are trying to develop adaptive algorithms for sequencing different kinds of instructional activities (e.g., using reinforcement learning). For instance, researchers designing intelligent tutoring systems have used ICAP as a general framework to devise adaptive scaffolding for cognitive engagement [4]. The authors found that, according to their models, adaptively choosing scaffolding activities is better than always choosing constructive scaffolds (i.e., the highest mode they could support based on ICAP). Our model can serve as a complementary theoretically grounded environment for comparing different data-driven approaches for adaptively sequencing activities or to gain insights into when algorithms should use different kinds of activities.
5.3 Bridging Between Conceptual and Procedural Learning
@@
While we claimed ICAP has primarily been studied in the context of conceptual learning, our model swings the pendulum to the opposite extreme of only focusing on procedural knowledge acquisition. In future work, we seek to develop a model that can account for both conceptual and procedural learning. This would be important for several reasons. First, we noted earlier that Assumption 2 (i.e., that knowledge construction is in terms of reinforcing edges in the knowledge graph) may be an oversimplification. In reality, we believe that constructively learning procedural skills leads to constructing conceptual knowledge. In Procedural ABICAP, conceptual knowledge is implicitly the edges between nodes. However, this is likely not sufficiently general. There may be a concept that connects more than two nodes, in which case constructing conceptual knowledge would transfer to several different procedural skills. The most natural way of incorporating this is by having both conceptual and procedural nodes in the knowledge graph. Second, it has been shown that learning procedural knowledge has a positive impact on learning corresponding conceptual knowledge and vice versa [12]. We could draw upon this body of literature to extend our model, which could give new insights into the relatively understudied bidirectional relationship between procedural and conceptual knowledge [12].

### v2

@@ Introduction
As shown in Figure 2, after enough steps, the ICAP hypothesis does indeed hold (i.e., $I>C>A>P$). However, what is more interesting is that this is not always the case. When learners have low prior knowledge, the ICAP hypothesis does not hold. In particular, until around step 35, we find that passive learning is consistently at least as good as (and typically better than) active and constructive learning. This is consistent with the worked example effect. However, after step 35, active and constructive learning gradually overtake passive learning, consistent with the expertise reversal effect. The reason for this is that early on, the increased cognitive load is too high to make active or constructive learning worthwhile, but as the learners learn more, the support of neighboring nodes (and the reinforcement of edges in the case of constructive learning) makes these kinds of learning activities more beneficial, despite the increased cognitive load. This shows why it could be important to take both cognitive load considerations and ICAP considerations into account to make accurate predictions about learning.
Discussion and Future Work
@@
Our model represents a more nuanced version of ICAP that is designed to account for empirical results on learning procedural knowledge. Our results show that by modeling cognitive load, the ICAP framework may be able to explain prior empirical literature beyond results on purely conceptual learning. Here we discuss three directions we could pursue in future work to truly capitalize on the benefits of Procedural ABICAP.
Replication Studies
@@
While our model can replicate some broad findings in the literature (e.g., the ICAP hypothesis and the expertise reversal effect), we note that these replication studies are still preliminary. In future work, we could do more precise replications of specific findings. For example, here we showed how conceptually our model could replicate the expertise reversal effect. However, we do not know if it would actually predict expertise reversal effects noticed in the literature. By trying to replicate specific studies that noticed such an effect (e.g., by mimicking their knowledge graphs and experimental setups), we may get better insights into the validity of our model and how it may need to be adjusted. Moreover, Procedural ABICAP could be used to replicate a variety of other results in the literature as well. For example, we could try to replicate results on studies that interleave worked examples and problem-solving tasks (?, ?).
Making Novel Predictions
@@
Although we have primarily focused on replicating prior results, the true power of agent-based modeling lies in making predictions about novel settings where there is little empirical literature. This could be used to generate hypotheses which could subsequently be verified with experimental studies (?, ?).
@@
Researchers who design interventions based on the ICAP framework often assume that learning gains from higher modes of engagement are higher (?, ?) but as we have shown with our model, this is not always the case. Using a computational model could help elucidate when we expect certain instructional interventions to be more effective than others. Of course, for the simulations to be accurate, it would help to have a concrete sense of the knowledge graph, the cognitive load for different modes of engagement, etc. If researchers have a lot of uncertainty about certain parameters (e.g., the relative cognitive load for different activities), they can use robustness tests to assess the sensitivity of the simulation results to the choice of those parameters. Moreover, models like Procedural ABICAP could also support researchers who are trying to develop adaptive algorithms for sequencing different kinds of instructional activities (e.g., using reinforcement learning). For instance, researchers designing intelligent tutoring systems have used ICAP as a general framework to devise adaptive scaffolding for cognitive engagement (?, ?). The authors found that, according to their models, adaptively choosing scaffolding activities is better than always choosing constructive scaffolds (i.e., the highest mode they could support based on ICAP). Our model can serve as a complementary theoretically grounded environment for comparing different data-driven approaches for adaptively sequencing activities or to gain insights into when algorithms should use different kinds of activities.
Bridging Between Conceptual and Procedural Learning
@@
While we claimed ICAP has primarily been studied in the context of conceptual learning, our model swings the pendulum to the opposite extreme of only focusing on procedural knowledge acquisition. We are currently working on developing a model that can account for both conceptual and procedural learning. This would be important for several reasons. First, we noted earlier that Assumption 2 (i.e., that knowledge construction is in terms of reinforcing edges in the knowledge graph) may be an oversimplification. In reality, we believe that constructively learning procedural skills leads to constructing conceptual knowledge. In Procedural ABICAP, conceptual knowledge is implicitly the edges between nodes. However, this is likely not sufficiently general. There may be a concept that connects more than two nodes, in which case constructing conceptual knowledge would transfer to several different procedural skills. The most natural way of incorporating this is by having both conceptual and procedural nodes in the knowledge graph. Second, it has been shown that learning procedural knowledge has a positive impact on learning corresponding conceptual knowledge and vice versa (?, ?). We could draw upon this body of literature to extend our model, which could in turn generate new insights into the relatively understudied bidirectional relationship between procedural and conceptual knowledge (?, ?).